### v1

# Dynamic linear regression models for forecasting time series with semi long memory errors

Abstract.
Dynamic linear regression models forecast the values of a time series based on a linear combination of a set of exogenous time series while incorporating a time series process for the error term. This error process is often assumed to follow an autoregressive integrated moving average (ARIMA) model, or seasonal variants thereof, which are unable to capture a long-range dependency structure of the error process. We propose a novel dynamic linear regression model that incorporates the long-range dependency feature of the errors. We demonstrate that the proposed error process may (i) have a significant impact on the posterior uncertainty of the estimated regression parameters and (ii) improve the model’s forecasting ability. We develop a Markov chain Monte Carlo method to fit general dynamic linear regression models based on a frequency domain approach that enables fast, asymptotically exact Bayesian inference for large datasets. We demonstrate that our approximate algorithm is faster than the traditional time domain approaches, such as the Kalman filter and the multivariate Gaussian likelihood, while retaining a high accuracy when approximating the posterior distribution. We illustrate the method in simulated examples and two energy forecasting applications.
Keywords: Dynamic linear regression, forecasting, Bayesian inference.
$\dagger$: School of Economics, University of New South Wales.
$\ddagger$: School of Mathematical and Physical Sciences, University of Technology Sydney.
§$\S$: Data Analytics Center for Resources and Environments (DARE)
1. Introduction
Forecasting time series data plays an important role in various fields, such as engineering, economics, and climate sciences. Forecasting a single output time series may be more accurate by using a linear combination of one or more exogenous time series that explains some of its historical variation (Hyndman and Athanasopoulos,, 2018). However, the linear combination of the exogenous time series often does not capture all the serial correlation present in the output time series, resulting in errors that are autocorrelated, i.e. serially correlated. Dynamic linear regression models (Pankratz,, 2012) provide a framework that relates the output time series to a linear combination of the exogenous time series and models the resulting error term as a time series process to account for serially correlated errors. Dynamic linear regression (DLR) models are a particular case of the more general class transfer function models (Box et al.,, 2015).
The standard dynamic linear regression model assumes that the error process, i.e. the time series process for the error terms, follows an autoregressive integrated moving average process (ARIMA). This approach allows the error process to have an autoregressive component and a moving average component for the white noise error. While this error process accommodates a wide range of stationary processes, it may require a large number of autoregressive components to accommodate error processes that show significant autocorrelation between distant time periods. To model the autocovariance of the error process parsimoniously, we use the autoregressive tempered fractionally integrated moving average (ARTFIMA) (Meerschaert et al.,, 2014; Sabzikar et al.,, 2019). The ARTFIMA class nests the well-known autoregressive fractionally integrated moving average (ARFIMA) (Granger and Joyeux,, 1980) model, which is useful for modelling time series with so-called long memory.
The ARTFIMA model has the following three advantages compared to the ARFIMA model, where the latter two are of particular importance for the frequency domain estimation approach for dynamic linear regression models proposed in this paper. First, the autocovariance function of the ARFIMA model decays at such a slow rate (power-law decay) that it is not absolutely summable, making it hard to analyse (Sabzikar et al.,, 2019). Such a process is referred to as having long memory. In contrast, the ARTFIMA model has a summable autocovariance function, which exhibits long-range dependence for a number of lags but eventually decays exponentially fast (i.e. much faster than power law). Such a process is referred to as having semi-long memory. Second, the spectral density of the ARFIMA process diverges as the frequency approaches zero. In practice, however, the empirical power spectrum (the estimate of the spectral density) is bounded for small frequencies. This is empirically illustrated in Meerschaert et al., (2014); Sabzikar et al., (2019), who show that the ARTFIMA process (whose spectral density does not diverge as the frequency approaches zero) fits the power spectrum better than the ARFIMA process for low frequencies; see also Figure 10 showing this carries over to the DLR case. This stylised fact is important for our purpose as we develop an estimation method based on a parametric Whittle likelihood (Whittle,, 1953), which is a frequency domain approximation of the time domain likelihood that represents the information in the data via the spectral density. Third, the Whittle approximation fails to hold for long memory processes, especially for small frequencies (Robinson,, 1995; Rousseau et al.,, 2012). Section 4.2 demonstrates this drawback of ARFIMA models via a simulation study and also shows that the ARTFIMA model overcomes this issue. We show that resorting to the time domain likelihood for estimating a dynamic linear regression with ARFIMA errors (Doornik and Ooms,, 2003) is computationally much more costly than our approach of a frequency domain likelihood with ARTFIMA errors. Moreover, in terms of prediction accuracy, our model is often superior, or at least on par with.
The model parameters in dynamic linear regression models consist of the regression coefficients that form the linear combination of the exogenous time series and the parameters of the error process. For the standard dynamic linear regression models with ARIMA errors that are normally distributed, efficient likelihood-based inference can be carried out by finding the finite-dimensional state space representation of the model and using the Kalman filter to integrate out the unobserved time-varying error terms. This may still be computationally costly with many time observations, especially in Bayesian inference, which typically requires many posterior samples for reliable inference. For each such sample, the Kalman filter needs to cycle through all observations, which can be prohibitively expensive, especially for large time series. The Whittle log-likelihood can be derived using large sample properties of the so-called periodogram data, which are formed via the discrete Fourier transform of the time domain data. The Whittle log-likelihood is directly a function of the regression coefficients and the error process parameters without needing to integrate out unobserved error terms. However, in dynamic linear regression models, the periodogram data becomes a function of the regression coefficients, which requires recomputing the discrete Fourier transform in every iteration. We show how this can be circumvented, and thus, our algorithm requires computing the discrete Fourier transform only once before running the sampling algorithm, thereby obtaining significant computational gains compared to the time domain log-likelihood based on the Kalman filter. Our frequency domain approach is also applicable to a dynamic linear regression model with ARTFIMA errors, where a finite-dimensional state space representation is not readily available.
To summarise, our article has two important contributions. First, we propose a frequency domain estimation approach for dynamic linear models that significantly outperforms estimation approaches based on the time domain likelihood in terms of computing time, especially when a finite-dimensional state space representation of the model is not available. Second, we utilise a semi-long memory process for the error process and show that it may provide more accurate forecasts compared to both the standard dynamic linear model and that of using a long memory process for the error term, especially for datasets with a complex long-range dependency. Moreover, we demonstrate that the semi-long memory process may significantly impact the posterior distribution of the regression parameter.
The rest of the article is organised as follows. Section 2 reviews existing dynamic linear models and presents our extension. Section 3 introduces the necessary frequency domain tools, outlines our estimation method, and validates its performance relative to the time domain likelihood. Section 5 presents applications for two real-world electricity demand datasets.
2. Dynamic linear regression models
2.1. Standard dynamic linear regression models
Let $\boldsymbol{X}_{t}=(X_{1t},\dots,X_{mt})^{\top}\in\mathbb{R}^{m}$ be a set of $m$ exogenous stationary time series observed at time $t$. A dynamic linear regression model models the single output time series $Y_{t}\in\mathbb{R}$ as a linear combination of the exogenous $\boldsymbol{X}_{t}$, i.e.
(2.1) $Y_{t}$ $=\boldsymbol{X}_{t}^{\top}\boldsymbol{\beta}+\eta_{t},$
where $\boldsymbol{\beta}\in\mathbb{R}^{m}$ is a vector of regression coefficients and $\eta_{t}\in\mathbb{R}$ is a zero-mean error process. The error process contains all unobserved factors that affect $Y_{t}$, and we assume that $\mathrm{E}(\eta_{t}|\boldsymbol{X}_{t})=0$, i.e. it is uncorrelated with each of the elements in $\boldsymbol{X}_{t}$ and hence $\boldsymbol{X}_{t}$ is exogenous.
When the error process $\eta_{t}$ does not show any temporal dependence, (2.1) is a standard linear regression (given it has a constant variance) and can easily be estimated using standard linear regression approaches (or modified versions thereof if heteroscedasticity is present). However, in many applications, the unobserved factors are time-varying, resulting in serially correlated errors $\eta_{t}$. The standard dynamic linear regression model assumes that $\eta_{t}$ is an autoregressive integrated moving average process, denoted ARIMA$(p,d,q)$ and defined as
(2.2) $\phi_{p}(B)\Delta^{d}\eta_{t}$ $=\psi_{q}(B)\varepsilon_{t},$
[…]
(2.3) $\phi_{p}(B)\phi_{P}^{\star}(B^{s})\Delta^{d}\Delta^{D}_{s}\eta_{t}=\psi_{q}(B)\psi_{Q}^{\star}(B^{s})\varepsilon_{t},$
where $\phi_{P}^{\star}(B^{s})=1-\sum^{P}_{i=1}\phi^{\star}_{i}B^{is}$ and $\psi_{Q}^{\star}(B^{s})=1+\sum^{Q}_{i=1}\psi^{\star}_{i}B^{is}$ are the seasonal autoregressive and seasonal moving average lag polynomials, and $\Delta^{D}_{s}=(1-B^{s})^{D}$, for $D=0,1,2,\dots$, is the seasonal differencing operator (Box et al.,, 2015).
To carry out inference in (2.1), one typically assumes that $\eta_{t}$ follows a normal distribution. The resulting log-likelihood is then a multivariate Gaussian distribution with a Toeplitz covariance matrix given that $\eta_{t}$ is stationary (Doornik and Ooms,, 2003). Inversion of such a matrix is usually performed via the Levinson-Durbin algorithm (Levinson,, 1946; Durbin,, 1960) in $\mathcal{O}(T^{2})$ operations. However, recent approaches, known as superfast Toeplitz algorithms, can solve the matrix system in $\mathcal{O}(T\log_{2}^{2}(T))$ operations, which becomes more efficient than the Levinson-Durbin algorithm when $T>256$ (Ammar and Gragg,, 1988). A method that scales even better is to find the finite-dimensional state space representation of the model in (2.1) with the error process following either (2.2) or (2.3), in which the resulting likelihood can be evaluated in $\mathcal{O}(T)$ using the Kalman filter.
2.2. Long memory dynamic linear regression models
In many applications, a pair of observations separated by a long time interval exhibit a non-negligible correlation. The standard dynamic linear regression model with the error process in (2.2) has an autocovariance function whose absolute value decays exponentially fast and can thus not capture this feature. A common approach to model time series with long memory is to consider so-called fractional differencing.
In a dynamic linear regression setting, Doornik and Ooms, (2004) propose to model the error process $\eta_{t}$ in (2.1) with an autoregressive fractionally integrated moving average process (Granger and Joyeux,, 1980; Hosking,, 1981), denoted ARFIMA$(p,d,q)$. This model is, for $d\notin\mathbb{Z}$,
(2.4) $\phi_{p}(B)\Delta^{d}\eta_{t}$ $=\psi_{q}(B)\varepsilon_{t},$
where $\phi_{p}(B)$ and $\psi_{q}(B)$ are defined as in (2.2). For $d\notin\mathbb{Z}$ the fractional differencing operator is defined via the fractional binomial theorem
$\Delta^{d}\eta_{t}$ $=(1-B)^{d}\eta_{t}$
$=\sum_{j=0}^{\infty}{\binom{d}{j}}(-B)^{j}\eta_{t}$
$=\sum_{j=0}^{\infty}(-1)^{j}\frac{\Gamma(1+d)}{\Gamma(1+d-j)j!}\eta_{t-j},$
[…]
Doornik and Ooms, (2003) show that maximum likelihood estimation for ARFIMA models can be carried out in $\mathcal{O}(T^{2})$ time, and this also applies for the model in (2.1) with the error process following (2.4). Chan and Palma, (1998) show that there is no finite-dimensional state space representation of an ARFIMA process. They propose an infinite-dimensional state space representation for which the Kalman filter can be computed in a finite number of steps equal to $T$; however, the resulting computation is $\mathcal{O}(T^{3})$.
2.3. Semi long memory dynamic linear regression models
The autoregressive tempered fractional integrated moving-average (ARTFIMA) (Meerschaert et al.,, 2014; Sabzikar et al.,, 2019) is an extension of the ARFIMA model that incorporates a tempering parameter $\lambda$. The model for the error process is then
(2.5) $\phi_{p}(B)\Delta^{d,\lambda}\eta_{t}$ $=\psi_{q}(B)\varepsilon_{t},$
where $\phi_{p}(B)$ and $\psi_{q}(B)$ are defined as in (2.2) and for $d\notin\mathbb{Z}$ and $\lambda>0$ the tempered fractional differencing operator is
$\Delta^{d,\lambda}\eta_{t}$ $=(1-\exp(-\lambda)B)^{d}\eta_{t}$
$=\sum_{j=0}^{\infty}{\binom{d}{j}}(-B)^{j}\eta_{t}$
(2.6) $=\sum_{j=0}^{\infty}(-1)^{j}\frac{\Gamma(1+d)}{\Gamma(1+d-j)j!}\exp(-\lambda j)\eta_{t-j}.$
Note that when $\lambda=0$, the ARTFIMA model becomes a stationary ARFIMA model if $-0.5<d<0.5$, provided the roots of the polynomial $\phi_{p}(z)$ are outside the unit circle in the complex plane. When $\lambda=0$ and $d$ is integer-valued, the ARTFIMA model becomes an ARIMA model. The ARTFIMA model is stationary for all $d\notin\mathbb{Z}$ and $\lambda>0$ provided that the root condition of the polynomial $\phi_{p}(z)$ is satisfied (Sabzikar et al.,, 2019). The ARTFIMA process has several advantages over the ARFIMA process, which are outlined in Section 1.
Similar to the ARFIMA model, a finite-dimensional state space representation that allows efficient Kalman filter computations is not readily available. Instead, the log-likelihood can be computed using a multivariate Gaussian distribution with a Toeplitz covariance matrix formed via the autocovariance function in Sabzikar et al., (2019, Theorem 2.5 (b)). However, a more computationally efficient approach is forming a frequency domain log-likelihood for the so-called periodogram data described in Section 3. The approach relies on the spectral density, also known as the power, of the process which we now describe.
Let $\omega$ denote the angular frequency. The spectral density $f(\omega)$ is the Fourier transform of the covariance function and shows how the variance of the time series is distributed among the frequencies $-\pi\leq\omega\leq\pi$. The spectral density when $\eta_{t}$ follows the ARTFIMA process in (2.5) is (Sabzikar et al.,, 2019, Theorem 2.5 (a))
(2.7) $f(\omega;\boldsymbol{\vartheta})$ $=\frac{\sigma_{\varepsilon}^{2}}{2\pi}\left|1-e^{-(\lambda+i\omega)}\right|^{-2d}\left|\frac{\psi_{q}(e^{-i\omega})}{\phi_{p}(e^{-i\omega})}\right|^{2},$
[…]
(2.8) $\phi_{p}(B)\phi_{P}^{\star}(B^{s})\Delta^{d,\lambda}\Delta^{D}_{s}\eta_{t}=\psi_{q}(B)\psi_{Q}^{\star}(B^{s})\varepsilon_{t},$
where all quantities have been previously defined, has the spectral density
(2.9) $f(\omega;\boldsymbol{\vartheta})$ $=\frac{\sigma_{\varepsilon}^{2}}{2\pi}\left|1-e^{-(\lambda+i\omega)}\right|^{-2d}\left|\frac{\psi_{q}(e^{-i\omega})\psi^{\star}_{Q}(e^{-is\omega})}{\phi_{p}(e^{-i\omega})\phi^{\star}_{P}(e^{-is\omega})}\right|^{2}.$
Figure 1 shows the spectral density for both the ARFIMA and ARTFIMA process for different values of the fractional differencing parameter $d$. An important difference that we have stressed is the limiting behaviour as $\omega\rightarrow 0$, where the spectral density of the ARFIMA process diverges when $d>0$, whereas that of the ARTFIMA process does not. When estimating the power for real data, it is often observed to be bounded. Hence the ARTFIMA model provides a better fit for small frequencies; see for example Figure 10.
Figure 1. Spectral densities of different ARFIMA and ARTFIMA models with $p=1$ and $q=0$. For both processes, the parameter $\phi=0.5$ and ARTFIMA $\lambda=0.045$.
3. Methodology
3.1. Frequency domain likelihood
The frequency domain approach to inference relies on the asymptotic properties of the frequency representation of the time series process in (2.1). Engle, (1974) shows how to rewrite (2.1) in terms of the periodograms of $Y_{t}$ and $\boldsymbol{X}_{t}$ and derive the ordinary least squares estimator of the transformed regression, which is also shown to be the best linear unbiased estimator. Our Bayesian approach requires a log-likelihood function, which can be derived using asymptotic properties of the periodogram data, which we now outline in some detail.
Let $Y_{t}$ and $\boldsymbol{X}_{t}$ be zero-mean processes. Suppose that $\boldsymbol{\beta}$ is known and define the pseudo data
$Z_{t}=Y_{t}-\boldsymbol{X}_{t}^{\top}\boldsymbol{\beta},\,\,t=1,\dots,T.$
The general case assumes that $\eta_{t}$ in (2.1) is the seasonal ARTFIMA process in (2.8) and hence the spectral density is given by (2.9). The spectral density of the ARTFIMA process in (2.7) is obtained using $Q=P=0$ in (2.9). The spectral density of the seasonal ARMA process is obtained by setting $\lambda=0$ and $d=0$ in (2.9), and for ARMA, in addition, $Q=P=0$.
Recall $\omega\in[-\pi,\pi]$ as the angular frequency and denote the natural Fourier frequencies as $\omega_{k}=\frac{2\pi k}{T}$ for $k\in\mathcal{K}$, where
$\mathcal{K}=\begin{cases}-\frac{T}{2},-\frac{T}{2}+1,\dots,\frac{T}{2}-1,\text{ if }T\text{ is even},\\ -\frac{(T-1)}{2},-\frac{(T-1)}{2}+1,\dots,\frac{(T-1)}{2},\text{ if }T\text{ is odd}.\end{cases}$
The frequency representation of the time series process $Z_{t}$ is obtained via its discrete Fourier transform (DFT), which is the complex-valued transform
(3.1) $J_{Z}(\omega_{k})$ $=\sum_{t=1}^{T}Z_{t}\exp(-i\omega_{k}t),$
which can be computed for all $T$ frequencies $\omega_{k}$, $k\in\mathcal{K}$, using $\mathcal{O}(T\log_{2}(T))$ operations via the fast Fourier transform (Cooley and Tukey,, 1965). Since the pseudo data depends on $\boldsymbol{\beta}$, the DFT in (3.1) needs to recomputed for each new sample in the Markov chain Monte Carlo (MCMC) algorithm. However, following Matsuda and Yajima, (2009),
$J_{Z}(\omega_{k})$ $=\sum_{t=1}^{T}(Y_{t}-\boldsymbol{X}_{t}^{\top}\boldsymbol{\beta})\exp(-i\omega_{k}t)$
$=\sum_{t=1}^{T}Y_{t}\exp(-i\omega_{k}t)-\left(\sum_{t=1}^{T}\boldsymbol{X}_{t}\exp(-i\omega_{k}t)\right)^{\top}\boldsymbol{\beta}$
$=J_{Y}(\omega_{k})-\boldsymbol{J}_{\boldsymbol{X}}(\omega_{k})^{\top}\boldsymbol{\beta},$
[…]
The DFT $J_{Z}(\omega_{k})$ is a weighted complex valued sum of the pseudo data. Peligrad and Wu, (2010, Theorem 2.1) show that under quite regular conditions,
(3.2) $\frac{1}{\sqrt{T}}(\Re(J_{Z}(\omega_{k})),\Im(J_{Z}(\omega_{k}))),\quad T\rightarrow\infty,$
where, respectively, $\Re(z)$ and $\Im(z)$ denote the real and imaginary parts of $z$, converge in distribution to a bivariate normal distribution with (asymptotically) independent components having expected value 0 and variance $\pi f(\omega_{k})$, with $f$ being the spectral density of $Z_{t}$. Moreover, they show that $\frac{1}{\sqrt{T}}J_{Z}(\omega_{k})$ are asymptotically independent for all $k\in\mathcal{K}$.
Define the periodogram
(3.3) $I_{Z}(\omega_{k})=\frac{1}{2\pi}\left|\frac{1}{\sqrt{T}}J_{Z}(\omega_{k})\right|^{2}.$
Then, for $k\in\mathcal{K}\setminus 0$, by (3.2),
(3.4) $\frac{I_{Z}(\omega_{k})}{f(\omega_{k})}$ $=\frac{1}{2}\left|\frac{1}{\sqrt{\pi f(\omega_{k})}}\frac{1}{\sqrt{T}}J_{Z}(\omega_{k})\right|^{2}\sim\frac{\chi^{2}(2)}{2},$
as $T\rightarrow\infty$, where $\chi^{2}(\nu)$ denotes the Chi-squared distribution with $\nu$ degrees of freedom. Note that $\nu=2$ follows from $\left|\cdot\right|^{2}$ in (3.4) being a sum of two squared (asymptotically) independent standard normal random variables when $k\in\mathcal{K}\setminus 0$. Since $\frac{\chi^{2}(2)}{2}$ is a standard exponential random variable, it follows that
(3.5) $I_{Z}(\omega_{k})$ $\sim\mathrm{Exp}(f(\omega_{k})),\,\,k\in\mathcal{K}\setminus 0,$
where $\mathrm{Exp}$ denotes an exponential random variable parameterised by its mean. Emphasising the dependence on the parameters, the log-density of (3.5) is
(3.6) $\log p(I_{Z}(\omega_{k};\boldsymbol{\beta})|\boldsymbol{\theta})$ $=-\log(f(\omega_{k};\boldsymbol{\vartheta}))-\frac{I_{Z}(\omega_{k};\boldsymbol{\beta})}{f(\omega_{k};\boldsymbol{\vartheta})},\,\,k\in\mathcal{K}\setminus 0,$
where $\boldsymbol{\theta}=(\boldsymbol{\vartheta},\boldsymbol{\beta})$ contains all unknown parameters. When $k=0$, $J_{Z}(0)=0$ since $J_{Y}(0)=\sum_{t=1}^{T}Y_{t}=0$ and $\boldsymbol{J}_{\boldsymbol{X}}(0)=\sum_{t=1}^{T}\boldsymbol{X}_{t}=\boldsymbol{0}$ for demeaned data.
The asymptotic distributions of the periodogram data underlie the idea of the so-called Whittle log-likelihood (Whittle,, 1953): Use the log-density of each periodogram observation in $I_{Z}(\omega_{k};\boldsymbol{\beta})$ (3.6) to form the log-likelihood of all periodograms, which becomes a sum due to the (asymptotic) independence. For a real-valued process $Z_{t}$, both the periodogram and spectral density are symmetric around the origin; hence, only non-negative frequencies are considered. The Whittle log-likelihood is obtained by adding (due to the asymptotic independence) the log-densities in (3.6) for all the positive frequencies (zeroth frequency excluded with demeaned data). The Whittle log-likelihood is, for odd $T$,
(3.7) $\ell_{W}(\boldsymbol{\theta})$ $=-\sum_{k=1}^{(T-1)/2}\left(\log(f(\omega_{k};\boldsymbol{\vartheta}))+\frac{I_{Z}(\omega_{k};\boldsymbol{\beta})}{f(\omega_{k};\boldsymbol{\vartheta})}\right),$
and the summation runs to $T/2-1$ instead if $T$ is even. It can be shown that the Whittle log-likelihood is an approximation of the Gaussian time domain log-likelihood (Guyon,, 1982; Kent and Mardia,, 1996).
Finally, we stress that a frequency domain approach for dynamic linear regression models with the Whittle log-likelihood is not appropriate when using the ARFIMA model because the Whittle approximation fails to hold for long memory processes, in particular for the small frequencies (Robinson,, 1995; Rousseau et al.,, 2012). Thus, to carry out inference in dynamic linear regression models with ARFIMA errors, one has to resort to the time domain likelihood, which is considerably slower to compute compared to a frequency domain approach that uses the ARTFIMA process; see Sections 4.1, 5.4, and 5.5.
3.2. Bayesian inference via Markov chain Monte Carlo
An important objective in time series is to learn $\boldsymbol{\theta}=(\boldsymbol{\vartheta},\boldsymbol{\beta})\subseteq\boldsymbol{\Theta}\subset\mathbb{R}^{\dim_{\boldsymbol{\vartheta}}+\dim_{\boldsymbol{\beta}}}$ given realisations of the time series processes $Y_{t}\subseteq\mathbb{R}$ and $\boldsymbol{X}_{t}\subseteq\mathbb{R}^{\dim_{\boldsymbol{X}}}$. Let $p(\boldsymbol{Z}|\boldsymbol{\theta})$ denote the likelihood function of $\boldsymbol{\theta}$ given the pseudo data $\boldsymbol{Z}=(Z_{1},\dots,Z_{T})$, which depends on the subset $\boldsymbol{\beta}$ of the parameter vector $\boldsymbol{\theta}$, but we suppress this dependence for notational simplicity. When the likelihood is obtained via the Whittle log-likelihood in (3.7),  $p(\boldsymbol{Z}|\boldsymbol{\theta})=\exp\left(\ell_{W}(\boldsymbol{\theta})\right)$. The likelihood function in this case is given the periodogram ordinates $I(\omega_{0}),\dots,I(\omega_{(T-1)/2})$, but since they are functions of $\boldsymbol{Z}$ we keep using the notation $p(\boldsymbol{Z}|\boldsymbol{\theta})$.
The cornerstone of Bayesian inference is the posterior distribution obtained via Bayes’ theorem,
[…]
where $p(\boldsymbol{\theta})$ is the prior distribution. MCMC methods are a class of iterative procedures to sample from (3.8), with the Metropolis-Hastings algorithm (Metropolis et al.,, 1953; Hastings,, 1970) arguably being the most used. The Metropolis-Hastings algorithm constructs a Markov chain $\left\{\boldsymbol{\theta}^{(j)}\right\}$ by starting at some initial value $\boldsymbol{\theta}^{(0)}$ and then, recursively, proposes a candidate draw $\boldsymbol{\theta}^{\prime}$ from a proposal density $q(\boldsymbol{\theta}|\boldsymbol{\theta}^{(j-1)})$ and sets $\boldsymbol{\theta}^{(j)}=\boldsymbol{\theta}^{\prime}$ with acceptance probability
(3.9) $\alpha_{\mathrm{MH}}=\min\left(1,\frac{p(\boldsymbol{Z}|\boldsymbol{\theta}^{\prime})p(\boldsymbol{\theta}^{\prime})/q(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{(j-1)})}{p(\boldsymbol{Z}|\boldsymbol{\theta}^{(j-1)})p(\boldsymbol{\theta}^{(j-1)})/q(\boldsymbol{\theta}^{(j-1)}|\boldsymbol{\theta}^{\prime})}\right).$
If a proposed draw is rejected, $\boldsymbol{\theta}^{(j)}=\boldsymbol{\theta}^{(j-1)}$. The acceptance probability in (3.9) is set so that, informally speaking, the Markov chain consists of samples from the posterior in (3.8) after a warm-up period referred to as the burn-in. Let $\left\{\boldsymbol{\theta}^{(j)}\right\}_{j=1,\dots,N}$ be the samples after discarding the burn-in. The proposal density is often taken to be a random walk, e.g.
$q(\boldsymbol{\theta}|\boldsymbol{\theta}^{(j-1)})=\mathcal{N}(\boldsymbol{\theta}|\boldsymbol{\theta}^{(j-1)},c\boldsymbol{\Sigma}_{\mathrm{prop}}),$
where $\mathcal{N}(\boldsymbol{x}|\boldsymbol{\mu},\boldsymbol{\Sigma})$ denotes the multivariate normal density of $\boldsymbol{x}$ with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$. Moreover, $\boldsymbol{\Sigma}_{\mathrm{prop}}$ is an approximation of the posterior covariance matrix and $c=2.38/\sqrt{\dim(\boldsymbol{\theta})}$ for optimality (Gelman et al.,, 1997).
If the algorithm rejects too often, the Markov chain becomes “sticky”, which causes inefficient estimates of posterior expectations. To explain how this inefficiency is measured, suppose $\theta$ is scalar-valued and consider estimating the expectation of a function $h$, i.e.
[…]
By the weak law of large numbers,
[…]
If the samples are independent, then the asymptotic variance of $\sqrt{N}\widehat{I}_{N}$ is $\sigma_{h}^{2}$, where $\sigma^{2}_{h}=\mathrm{Var}(h(\theta))$. However, MCMC results in correlated samples and then the asymptotic variance of $\sqrt{N}\widehat{I}_{N}$ is $\sigma_{h}^{2}(1+\sum_{i=1}^{\infty}\rho_{i})$, where $\rho_{i}$ is the autocorrelation at the $i$th lag the Markov chain. The term $(1+\sum_{i=1}^{\infty}\rho_{i})$ is called the inefficiency factor. It measures how much the asymptotic variance of the estimate of a posterior expectation is inflated compared to an ideal sampler that produces independent draws. The effective sample size (ESS) of the resulting Markov chain is defined as $\text{ESS}=N/(1+\sum_{i=1}^{\infty}\rho_{i})$. Thus, the effective sample size is an estimate of the number of samples equivalent to independent samples the sampler is generating.
3.3. Parameterisations and prior distributions
The autoregressive parameters $\phi_{i}$ are reparameterised in terms of partial autocorrelations $-1\leq\widetilde{\phi}_{i}\leq 1$, for $i=1,\dots,p$, following Barndorff-Nielsen and Schou, (1973) to ensure the process is always stationary. For the moving average parameters $\psi_{j}$, for $j=1,\dots,q$, the same reparameterisation in terms of $\widetilde{\psi}_{j}$ ensures the process is always invertible. To make the variance parameter $\sigma^{2}_{\varepsilon}$ and the tempering parameter $\lambda$ unrestricted, we reparameterise using log-transformations. For the ARTFIMA process, the fractional parameter $d$ ($\notin\mathbb{Z}$) is unrestricted (since stationarity does not depend on $d$). For ARFIMA models, on the other hand, $-0.5<d<0.5$ is a necessary condition for stationarity. We thus reparameterise $d$ to an unrestricted parameter $\widetilde{d}$ by a scaled Fisher transformation, i.e. $\widetilde{d}=\mathrm{arctanh}(2d)$.
When proposing a sample from the multivariate normal proposal as described in Section 3.2, a prior on the interval $[-1,1]$ for each autoregressive and moving average parameter ensures $|\widetilde{\phi}_{i}|\leq 1$ and $|\widetilde{\psi}_{j}|\leq 1$ for the $i$ and $j$ above. We set independent (within and between) uniform priors $\boldsymbol{\phi}\sim\mathrm{Unif}(-1,1)^{p}$ and $\boldsymbol{\psi}\sim\mathrm{Unif}(-1,1)^{q}$. The priors for $\log(\sigma^{2}_{\varepsilon})$ and $\log(\lambda)$ are $\mathcal{N}(0,100)$. For the ARTFIMA process $d\sim\mathcal{N}(0,1)$, while for ARFIMA models $\widetilde{d}\sim\mathcal{N}(0,1)$ which implies a weakly informative on $(-0.5,0.5)$ for $d$. Finally, $\beta_{k}$, for $k=1,\dots,m$, are assigned independent $\mathcal{N}(0,100)$ priors. We have verified that our results are not influenced by the choice of prior (the datasets studied are very large).
4. Simulated examples
4.1. Approximating the time domain likelihood
We compare posterior inference via the Whittle likelihood to those of using the exact Gaussian and Kalman filter likelihoods under the same weakly informative priors for dynamic linear regression models on simulated data. By the exact Gaussian likelihood, we mean evaluating a multivariate normal density where the covariance matrix is computed from the autocovariance function. The Kalman filter likelihood, on the other hand, performs the Kalman filter recursions to compute the likelihood. Unlike the Gaussian likelihood approach, the Kalman filter avoids inverting a large covariance matrix. Note that both these methods use the same (exact) time domain likelihood, however, evaluated in a different manner. This example is restricted to an ARMA process as the ARFIMA/ARTFIMA likelihoods cannot be efficiently evaluated using the Kalman filter.
Let $y_{t}=\beta_{1}x_{t}+\eta_{t}$ where $\eta_{t}$ is an ARMA(3,1) model with $T=10{,}001$. We set $\beta_{1}=3$ and the reparameterised vector
$\boldsymbol{\vartheta}=\left(\widetilde{\phi}_{1},\widetilde{\phi}_{2},\widetilde{\phi}_{3},\widetilde{\psi}_{1},\log(\sigma_{\varepsilon}^{2})\right)^{\top}=\left(0.4,-0.2,0.1,-0.2,\log(2)\right)^{\top}.$
The exogenous predictor $x_{t}$ is assumed to be known and is simulated from an ARMA(2,2) process. MCMC is used to sample the three posteriors via $10{,}000$ iterations of the random walk Metropolis-Hastings algorithm. Figure 2 plots the kernel density estimates of the marginal posteriors for each (untransformed) parameter and shows that the three posteriors are almost identical. Hence, the $5{,}000$ ($(T-1)/2$) periodogram ordinates are sufficient to give a very accurate approximation of the true time domain likelihood in this case.
Figure 2. Kernel density estimates of the marginal Whittle, Kalman filter and Gaussian posteriors for a DLR model with $\eta_{t}\sim$ ARMA(3,1).
Figure 3 reports the effective sample sizes. The figure shows that the Gaussian likelihood has a $\approx 25\%$ higher ESS for the $\beta_{1}$ parameter compared to the Whittle likelihood and the Kalman filter. The Whittle approach has a lower ESS for $\sigma^{2}$ compared to its time domain counterparts. The AR and MA parameters are comparable for all three posteriors, with the Whittle posterior having slightly higher ESS for all of them.
Figure 3. Effective sample sizes for $10{,}000$ MCMC iterations for each parameter in a DLR model with $\eta_{t}\sim\mathrm{ARMA}(3,1)$. The results are shown for the three methods in the legend.
Table 1 presents the run times (in seconds) of each method. Our proposed Whittle approach is roughly 8 times faster than the Gaussian likelihood and 94 times faster than the Kalman filter. The Gaussian likelihood uses the SuperGauss package in R (Ling and Lysy,, 2017), which scales as $\mathcal{O}(T\log_{2}^{2}(T))$ from the superfast Toeplitz algorithm proposed in Ammar and Gragg, (1988). Despite the Kalman filtering being performed in $\mathcal{O}(T)$ operations, the algorithm sweeps through all observations recursively for each new MCMC proposal and is thus time-consuming, especially in the Python implementation used in this paper. The Whittle log-likelihood (3.7), on the other hand, is a simple sum that only requires computing $I_{Z}(\omega)$ in (3.3) and the relevant spectral density, and is thus much faster than the Kalman filter.
Run time Whittle Gaussian Kalman filter
total (s) 10.0 80.54 942.58
Table 1. Computation time for sampling $10{,}000$ MCMC draws for each method.
Figure 4. Quantile-quantile (QQ) plots of the ratio $I_{Z}(\omega_{k})/f(\omega_{k})$ for simulated ARTFIMA vs ARFIMA models for the three lowest positive frequencies. The top row, middle row, and bottom row, correspond to $T=1001$, $T=10001$, and $T=20001$, respectively.
4.2. The periodogram distribution for low frequencies
This section empirically verifies the periodogram distribution for small frequencies for the ARTFIMA process. We also illustrate how this fails for the ARFIMA process. The latter has been investigated theoretically in Robinson, (1995); Rousseau et al., (2012) and empirically in Meerschaert et al., (2014); Sabzikar et al., (2019).
We test both models without the regression coefficient, i.e. $\boldsymbol{\beta}=\boldsymbol{0}$, and using
  1. (M1)
    ARTFIMA$(2,d,\lambda,0)$ with $(\phi_{1},\phi_{2},d,\lambda,\sigma^{2})=(0.742,0.227,2.139,0.616,1)$,
  2. (M2)
    ARFIMA$(2,d,0)$ with $(\phi_{1},\phi_{2},d,\sigma^{2})=(1.466,-0.525,0.493,1)$.
These parameter values are chosen to match those obtained by the real data application in Section 5.5. It follows from (3.5) that the ratio of the periodogram and its spectrum follows the distribution
(4.1) $\frac{I_{Z}(\omega_{k})}{f(\omega_{k})}\sim\frac{\chi^{2}(2)}{2},$
for $\omega_{k}\in\Omega$ as $T\rightarrow\infty$.
We verify (4.1) by simulating $10{,}000$ realisations for $T=1{,}001,10{,}001,20{,}001$ and keeping the three smallest positive Fourier frequencies $\omega_{k}$ (excluding zero). Figure 4 displays the quantile-quantile (QQ) plots of the simulated quantiles vs theoretical quantiles of (4.1) for all three cases of $T$. The left and right panels display the plots for the ARTFIMA and ARFIMA models, respectively. The upper, middle, and bottom rows correspond to $T=1{,}001,10{,}001,20{,}001$, respectively. The left panel of Figure 4 shows that for the three lowest frequencies, the empirical quantiles of the periodogram ratios are almost identical to their theoretical counterpart for all $T$. This confirms the suitability of (4.1) for the small frequencies in the ARTFIMA model case. In contrast, the right panel shows that (4.1) is not suitable for the small frequencies in the ARFIMA model case, and that an increasing $T$ does not make (4.1) more suitable.
The violation of (4.1) is partly attributed to the fact that the spectrum of the ARFIMA model behaves as a power law and diverges as $\omega\rightarrow 0$. Sabzikar et al., (2019) and Meerschaert et al., (2014) show for applications such as hydrology, finance, and climatology, the ARFIMA spectral density provides a poor fit of the periodogram at the low frequencies. We show the same issue in a dynamic linear regression model with ARFIMA errors; see Figure 10 in Section 5.4. A pragmatic solution is to use a low-frequency cutoff in the estimation of ARFIMA models in the frequency domain, however, this incurs a substantial loss of information in the estimation of the long-memory parameter $d$. Section 5 thus uses the exact Gaussian likelihood for dynamic linear models with ARFIMA errors, which corresponds to the exact maximum likelihood approach in Doornik and Ooms, (2004). The resulting computation is much slower than a frequency domain approach and stresses the importance of our contribution, namely: A computationally fast dynamic linear model that accounts for long-range dependence in the error process.
5. Applications
5.1. Overview and objectives
This section analyses electricity demand as the response variable conditional on relevant exogenous variables for two real-world datasets. The two datasets contain data for New England, United States of America (Example 1), and Victoria, Australia (Example 2), presented in Sections 5.4 and 5.5, respectively. Forecasting Example 2 is considerably more challenging than Example 1 and we show that accounting for long-range dependence is essential to obtain good forecast accuracy for moderately long horizons. In contrast, long-range dependency does not improve the forecasts in Example 1, however, we show it has a significant impact on the posterior distribution of the regression parameters $\boldsymbol{\beta}$.
We compare the estimation times and forecasting abilities of models with three types of error processes: ARTFIMA, ARFIMA, and ARMA. Section 5.2 discusses the choice of model settings and Section 5.3 presents the metrics used for the forecasting evaluation. Additionally, Example 1 illustrates the failure of a dynamic linear model with an ARFIMA error process to fit the observed power spectrum for the low frequencies.
5.2. Model settings
Given an error process (ARTFIMA, ARFIMA, or ARMA) we select the orders of $p$ and $q$ via the smallest Bayesian information criterion (BIC) value (Schwarz,, 1978). We restrict $0\leq p\leq 3$ and $0\leq q\leq 2$ for Example 1, and $0\leq p\leq 2$ and $0\leq q\leq 1$ for Example 2. The reason for smaller lag orders for the second example is that this example includes additional seasonal lags. After a given order $p,q$ for an error process is determined, we compare its forecasting ability to the two remaining error processes using the same $p,q$. We do this for all three error processes, so that each one has the opportunity to forecast using its optimal lag orders (the lag order may be the same for different error processes).
It is well known that the log-likelihood function for ARMA models may exhibit multimodality, and this is also true for ARTFIMA and ARFIMA. To avoid finding a poor local mode, we use global optimisation via Basin-hopping (Li and Scheraga,, 1987) to find the maximum likelihood estimate when computing the BIC.
5.3. Forecasting evaluation settings
For a given $p$ and $q$ determined as described in Section 5.2, the error processes are compared with respect to their forecasting abilities in a time series cross-validation setting. We use $k=100$ out-of-sample observations with different forecast horizons, $h=1,\dots,15$. The exogenous variables $\boldsymbol{X}_{t}$ are not modeled, i.e. their observed values are used for any forecasting horizon. A model to forecast future values of $\boldsymbol{X}_{t}$ is needed in practice, however, this simplifying assumption is made for all cases and does not affect the relative performance.
We evaluate both point and distributional forecasts. The optimal $h$-step ahead point forecast under a squared loss function is the conditional expectation
(5.1) $\mu_{y^{(h)}_{T+i}|\mathcal{I}_{T+i-h}}=\mathrm{E}(y_{T+i}|\mathcal{I}_{T-h-i}),$
where $\mathcal{I}_{b}$ denotes the information set at $t=b$, i.e. all the variables known up to and including time $t=b$. We evaluate the accuracy of the point forecast via the (out-of-sample) root mean squared error (RMSE)
$\text{RMSE}^{(h)}=\sqrt{\sum^{k}_{i=1}\frac{\left(y_{T+i}-\widehat{y}^{(h)}_{T+i}\right)^{2}}{k}},$
where $\widehat{y}^{(h)}_{T+i}$ is an estimate of (5.1) obtained using the maximum a posteriori (MAP) estimate for each model. To assess distributional forecasts, we obtain the $h$-step ahead predictive posterior density
[…]
Given the abundance of training data, i.e. a very large $T$, the posterior distribution $p(\boldsymbol{\theta}|y_{1:T-h-i})$ remains fairly unchanged for varying values $h\leq 15$ and $i\leq k=100$. Thus, for computational convenience, we avoid reestimating the model and sample $10{,}000$ MCMC draws from the posterior conditional on the full training data. The same posterior samples are then used to compute (5.2), i.e. only the conditional likelihood part of the integrand changes. The distributional forecasts are assessed via the continuous rank probability score (CRPS), which is defined as
[…]
where $F$ is the cumulative distribution function of the forecast and $\mathbbm{1}$ is the indicator function, taking the value one if the argument is true or zero otherwise (Matheson and Winkler,, 1976). When the distributional forecasts $F$ and the data distribution are equal, the CRPS achieves its minimum. We use 300 samples from the posterior predictive distribution in (5.2) to estimate the empirical cumulative distribution function $\widehat{F}(\widetilde{y}_{T+i})$ for all, $i=1,\dots,k=100$. The estimated CRPS is then summed over all time points to obtain $\widehat{\text{CRPS}}^{(h)}$, for $h=1,\dots,15$.
5.4. Example 1: New England electricity demand
The New England data consists of hourly electricity demand and temperature data for the states of Maine and Vermont from 1st March 2003 to April 30th, 2017, with $T=124{,}200$ observations. The response, Maine electricity demand, $y_{t}$, is modelled as a linear combination of lagged exogenous variables, Maine temperature $x_{1,t-1}$ and Vermont (neighbouring state) electricity demand $x_{2,t-1}$. Strong multi-seasonal and trend components exist within each variable. The mstl function in the R package Forecast (Hyndman and Khandakar,, 2008) is used to remove these effects. Figure 5 depicts each variable after removing the multi-seasonal and trend components and applying a Box-Cox transformation (Box and Cox,, 1964). As seen from the figure, the response exhibits long memory, which makes it a suitable candidate for an ARFIMA or ARTFIMA error process.
Figure 5. Maine and Vermont electricity demand (megawatts) and temperature (degrees Celsius) time series data after removing multi-seasonal and trend components and applying a Box-Cox transformation, alongside their corresponding autocorrelation plots. The measurement units are for the raw data.
The search over the model space described in Section 5.2 resulted in the following lag-orders for each error process: ARTFIMA$(1,d,\lambda,2)$, ARFIMA$(3,d,1)$, ARMA$(3,1)$. We thus have two cases: $p=1,q=2$ and $p=3,q=1$.
Table 2 reports the BIC values and average computation time for a log-likelihood evaluation for all three error processes for the case $p=1,q=2$. The table shows that ARFIMA has the smallest BIC value, however, it is roughly 40 times slower than both ARFIMA and ARMA since it uses the Gaussian log-likelihood. ARTFIMA has the second smallest BIC value and is only marginally slower than the ARMA model.
${\text{ARTFIMA}(1,d,\lambda,2)}$ ${\displaystyle\text{ARFIMA}(1,d,2)}$ ${\displaystyle\text{ARMA}(1,2)}$
BIC -383273.75 -383578.45 -382749.34
Time (ms) 11.3 403.0 10.5
Table 2. BIC values and the average time for one log-likelihood evaluation for each dynamic linear regression model for New England electricity demand.
Figure 6 plots the RMSE and CRPS for each error process for different forecast horizons $h$. The RMSE for all models are similar for small $h$; however, $h>6$ shows a lower RMSE score for the ARMA and ARTFIMA models compared to ARFIMA, with the ARMA model performing the best for $h>7$. The CRPS is similar for all models, with the ARFIMA model having the overall lowest and ARMA performing worse over the forecast horizon. Figure 7 plots the predictions and prediction intervals for the three error processes when predicting the electricity demand one and six hours ahead, i.e. $h=1,6$. All three error processes perform accurate predictions, with the prediction intervals getting wider for larger horizons as expected. To summarise the results, we conclude that ARMA and ARTFIMA are computationally much more efficient than ARFIMA. ARMA performs surprisingly well for forecasting, which indicates that the long-range dependency feature of the ARTFIMA and ARFIMA models does not strongly influence the forecasts in this case. However, inspecting the posterior density of $\beta$ for the different error processes in Figure 8 reveals a significant difference.
Figure 6. Root mean square error and the continuous rank probability scores for $h$-step ahead forecasts for New England electricity demand for $p=1,q=2$.
Figure 7. Prediction and prediction intervals for one hour ahead ($h=1$, top panel) and ($h=6$, bottom panel) predictions for three error processes (columns). The results are shown for New England electricity demand for $p=1,q=2$.
Figure 8. Kernel density estimator of the posterior density of $\boldsymbol{\beta}$. using $10{,}000$ MCMC samples. The results are shown for the New England data with $p=1,q=2$ and three error processes, ARTFIMA, ARFIMA, and ARMA.
Next, we consider the case of $p=3,q=1$. The MCMC sampler had problems converging for the ARTFIMA model; thus, the CRPS was computed at the MAP instead of using the full posterior predictive distribution. Table 3 displays the BIC values for each model (the relative computing times are similar to the $p=1$ $q=2$ case and therefore omitted). ARTFIMA has the smallest BIC value, followed by ARFIMA and ARMA. However, in terms of RMSE forecasts in Figure 9, ARMA performs the best. In terms of CRPS, both ARTFIMA and ARFIMA perform better than ARMA.
${\text{ARTFIMA}(3,d,\lambda,1)}$ ${\displaystyle\text{ARFIMA}(3,d,1)}$ ${\displaystyle\text{ARMA}(3,1)}$
BIC -383696.40 -383608.18 -382959.42
Table 3. BIC values for each dynamic regression model for New England electricity demand.
Figure 9. Root mean square error and the continuous rank probability scores for $h$-step ahead forecasts for New England electricity demand for $p=3,q=1$.
Figure 10 plots the fitted spectral densities (log-log scale) of the dynamic regression models with ARTFIMA and ARFIMA errors together with their corresponding periodogram data. Due to the dependence on $\boldsymbol{\beta}$, the two periodograms are different for every frequency $\omega$, but are almost identical at the lower frequencies. Consistent with the results of Sabzikar et al., (2019), the ARTFIMA spectrum provides a better overall fit to the periodogram data, in particular for small frequencies.
Figure 10. Spectral densities at the MAP and their respective periodograms. The results are shown for the New England data with $p=3,q=1$. The ARTFIMA spectrum (black line) with its periodogram (grey circles) and the fitted ARFIMA (orange line) with its corresponding periodogram (orange dots).
5.5. Example 2: Victorian electricity demand
The second application we consider is the half-hourly electricity demand for Victoria, Australia. The data consists of $T=52{,}608$ observations of electricity demand (in megawatts) and Melbourne’s temperature (degrees Celsius) between the 1st January 2012 - 31st December 2014. The response $y_{t}$, Victorian electricity demand, is modelled as a linear combination of the lagged exogenous variable Melbourne temperature $x_{t-1}$. Strong multi-seasonal and trend components exist within each variable and were removed via the mstl function. Figure 11 depicts each variable after removing the multi-seasonal and trend components and applying a Box-Cox transformation.
Figure 11. Victoria, Australia electricity demand (in megawatts) and temperature (degrees Celsius) time series data after removing multi-seasonal and trend components and applying a Box-Cox transformation, alongside their corresponding autocorrelation plots. The measurement units are for the raw data.
Despite removing seasonality, as seen from Figure 11, all seasonality cannot be removed and still exists at multiples of the 48th lag (1 day is 48 half-hours). This is also present in the periodogram, see Figure 14, which peaks at the corresponding frequencies.
All models achieve their smallest BIC for $p=2,q=1$ following the procedure described in Section 5.2. In addition, we include one seasonal MA term $Q=1$ with $s=48$ for ARMA and ARTFIMA and denote this parameter $\Psi^{\star}_{1}$ as in (2.3). We do not include a seasonal ARFIMA model as the Gaussian likelihood is expensive to compute and our paper has overwhelming evidence that ARTFIMA captures sufficient long-range dependence. Table 4 displays the BIC and the average time for one log-likelihood evaluation for the five models. The seasonal models have substantially lower BIC values than the non-seasonal models. The seasonal ARTFIMA error process gives the smallest BIC, followed by the seasonal ARMA. The table also shows that significant computational speedups are gained for ARTFIMA and ARMA models due to the frequency domain approach.
${\scriptstyle\text{ARTFIMA}(2,d,\lambda,1)(0,0,1)_{48}}$ $\scriptstyle{\text{ARTFIMA}(2,d,\lambda,1)}$ ${\scriptstyle\text{ARFIMA}(2,d,1)}$ ${\scriptstyle\text{ARMA}(2,1)}$ ${\scriptstyle\text{ARMA}(2,1)(0,0,1)_{48}}$
BIC -525257.64 -512383.65 -512102.13 -511832.00 -520631.12
Time 35.1ms 3.96ms 203.0ms 3.79ms 35.1ms
Table 4. BIC values for each dynamic regression model for Victorian electricity demand.
This dataset is particularly challenging and the MAP estimates were on the border of the stationarity or invertibility region (when transformed back from the partial autocorrelation parameterisation) for all models. As a consequence, MCMC sampling was difficult to carry out with reasonable efficiency. Thus, the CRPS (that requires posterior sampling) was only computed using the MAP estimate. Figure 12 reports the forecast RMSE and CRPS for the five models for all horizons $h=1,\dots,15$. The seasonal models have the lowest RMSE and CRPS compared to the non-seasonal models. The seasonal ARTFIMA model has the best forecasting ability in terms of both point estimation (RMSE) and distributional forecasts (CRPS). Figure 13 plots the predictions and prediction intervals for the seasonal models when predicting the electricity demand one and six hours ahead, i.e. $h=2,12$. While the one hour ahead prediction is accurate for both models (albeit slightly more accurate for ARTFIMA), the seasonal ARTFIMA clearly performs better for the six hours ahead forecast.
Figure 12. Root mean square error and the continuous rank probability score of the $h$-step ahead forecasts for Victorian electricity data for $p=2,q=1$
Figure 13. Prediction and prediction intervals for one hour ahead ($h=2$, top panel) and ($h=12$, bottom panel) predictions for two seasonal error processes (columns). The results are shown for Victorian electricity data for $p=2,q=1$.
Finally, Figure 14 further shows the appropriateness of the seasonal ARTFIMA model by displaying a log-log plot of the periodogram and spectral density at the MAP. It is evident that the seasonal ARTFIMA model fits the lower frequencies $\text{log}(\omega)<-4$ well, and also captures the seasonal features of the periodogram.
Figure 14. Main plot: Log frequency vs log power (log-log) plot of the periodogram and the spectral density function at the MAP for DLR with ARTFIMA errors. Zoomed-in plot: Similar figure but displaying moderate frequencies on a linear frequency scale.
6. Conclusion and future research
We proposed a dynamic linear regression models with an ARTFIMA error process that accounts for long-range dependence. We developed a frequency domain estimation method that utilises the computationally efficient FFT. The frequency domain method uses a likelihood based on asymptotic properties of the periodogram, and we demonstrated empirically that the resulting inference matches well with time domain likelihood methods. This result was expected as the case without an exogenous variable has strong theoretical results regarding the approximation error vanishing asymptotically.
We showed empirically that the well-known phenomena of the spectral density of the ARFIMA model not fitting the periodogram data well for small frequencies carries over to the dynamic linear regression case. As such, we argue that dynamic linear regression models with ARFIMA errors are not well suited for frequency domain methods, and their estimation needs time domain methods which are computationally expensive. In contrast, we showed that the dynamic linear regression models with ARTFIMA errors fit the low-frequency spectrum well and are thus a computationally fast alternative to model long-range dependence.
Our proposed model demonstrated superior forecasting performance when the data exhibited complex long-range dependency and was marginally worse if not on par with the ARMA error process for less complex data. This is expected as our model nests the dynamic linear regression model with ARMA errors. We did not see any signs of overfitting when using an ARTFIMA error process instead of an ARMA for the simpler dataset. In addition, we also observed that the posterior distribution of the regression parameter could be significantly affected by not accounting for long-range dependency in the error process. In summary, we believe dynamic linear regression models with ARTFIMA errors are robust methods that should be routinely included in a forecaster’s toolkit.
Future research will extend our approach to modelling a multivariate response as in Villani et al., (2022). For small to moderate datasets, the Whittle log-likelihood is known to be severely biased. Exploring recent advances such as the debiased Whittle likelihood (Sykulski et al.,, 2019) for dynamic linear regression models can potentially enable our method for small to moderately large datasets. Finally, the structure of our model can utilise subsampling MCMC methods (Quiroz et al.,, 2019, 2021) in the spectral domain (Salomone et al.,, 2020).